OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. \Show work please! Thanks!!! :D 3x + 4y = -2 and 2x - 3y = 10

OpenStudy (anonymous):

i think i answered this with steps in previous post

OpenStudy (saifoo.khan):

satallite to the resuce!

OpenStudy (anonymous):

i am sure i did.

OpenStudy (anonymous):

no u didnt now answer it por favor!

OpenStudy (saifoo.khan):

i got your answer!

OpenStudy (anonymous):

yay!!!!!! post it then, saif!

OpenStudy (saifoo.khan):

alright

OpenStudy (saifoo.khan):

2, -2

OpenStudy (katrinakaif):

So get an equation that solves for x. 2x - 3y = 10 Add 3y on the other side = 2x = 10 + 3x Divide it by 2 x = 10+ 3y /2 Now plug it into the other equation: 3x + 4y = -2 Subsitute in: 3 (10 + 3y/2) + 4y = -2 Distribute 3 gives us : 30/2 + 9y/2 + 4y = -2 Simplify 15 + 9y/2 + 4y = -2 Get similar denomators for 9y/2 and 4y/2 giving 15 + 9y/2 + 8y/2 = -2 Bring 15 onto the other side by substracting +9y/2 + 8y/2 = -17 +17y/2 = -17 Times by 2 to get rid of denominator 17y = -17 * 2 17y = -34 y = -2

OpenStudy (katrinakaif):

To solve for x plug it into orginal equation (y = -2) 3x + 4y = -2 3x + 4(-2) = -2 3x -8 = -2 Add 8 to the other side 3x = -2 + 8 3x = 6 x = 2